$SPAD/src/input roman.input

Timothy Daly December 3, 2016

Abstract

Contents

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool roman.output
)set message test on
)set message auto off
)clear all
--S 1 of 10
f := operator 'f
--R
--R
--R   (1)  f
--R                                                          Type: BasicOperator
--E 1
--S 2 of 10
D(f x,x,7)
--R
--R
--R         (vii)
--R   (2)  f     (x)
--R
--R                                                    Type: Expression(Integer)
--E 2
--S 3 of 10
a := roman(1978 - 1965)
--R
--R
--R   (3)  XIII
--R                                                           Type: RomanNumeral
--E 3
--S 4 of 10
x : UTS(ROMAN,'x,0) := x
--R
--R
--R   (4)  x

--R                               Type: UnivariateTaylorSeries(RomanNumeral,x,0)
--E 4
--S 5 of 10
recip(1 - x - x**2)
--R
--R
--R   (5)
--R                 2        3      4         5         6        7          8
--R     I + x + II x  + III x  + V x  + VIII x  + XIII x  + XXI x  + XXXIV x
--R   +
--R         9           10      11
--R     LV x  + LXXXIX x   + O(x  )
--R                    Type: Union(UnivariateTaylorSeries(RomanNumeral,x,0),...)
--E 5
--S 6 of 10
m : MATRIX FRAC ROMAN
--R
--R                                                                   Type: Void
--E 6
--S 7 of 10
m := matrix [[1/(i + j) for i in 1..3] for j in 1..3]
--R
--R
--R        + I    I    I+
--R        |--   ---  --|
--R        |II   III  IV|
--R        |            |
--R        | I    I   I |
--R   (7)  |---  --   - |
--R        |III  IV   V |
--R        |            |
--R        | I    I    I|
--R        |--    -   --|
--R        +IV    V   VI+
--R                                         Type: Matrix(Fraction(RomanNumeral))
--E 7
--S 8 of 10
inverse m
--R
--R
--R        +LXXII   - CCXL    CLXXX +
--R        |                        |
--R   (8)  |- CCXL    CM     - DCCXX|
--R        |                        |
--R        +CLXXX   - DCCXX    DC   +
--R                              Type: Union(Matrix(Fraction(RomanNumeral)),...)

--E 8
--S 9 of 10
y := factorial 10
--R
--R
--R   (9)  3628800
--R                                                        Type: PositiveInteger
--E 9
--S 10 of 10
roman y
--R
--R
--R   (10)
--R  ((((I))))((((I))))((((I)))) (((I)))(((I)))(((I)))(((I)))(((I)))(((I))) ((I))(
--R  (I)) MMMMMMMMDCCC
--R                                                           Type: RomanNumeral
--E 10
)spool
)lisp (bye)

———-

References^{[1] nothing}

5